please help me,,i'm so very confuse in this problem....
The measure of angle A is 9 more than twice the measure of angle B.If angle A and angle B are supplementary angles,what is the measure of angle A?

Answer by sudhanshu_kmr(1152)


angle B = x
angle A = 2x + 9
both are supplementary angles so,
x +2x+9 = 180
=> 3x = 180-9
=> 3x = 171
x = 171/3 = 57
angle B= 57
angle A = 123